OpenStudy (anonymous):

Would 2/5x - 7/20x = 3 be 60?

OpenStudy (anonymous):

Yes it works, when you put 60 in for x it will equal to 3.

OpenStudy (anonymous):

Yes, because 2/5 x - 7/20 x = 1/20 x So if you have 1/20 x = 3, cross-multiply, you get x = 60